# Mean-field Underdamped Langevin Dynamics and its Space-Time Discretization

### 26 Dec 2023 v1

@@ 4.1 Training mean-field neural networks
The objectives (14) satisfy Assumption 1-4 for specific common choices $l$, $h$ and $r$ described in several works  (Nitanda et al., 2022; Chen et al., 2022, 2023; Suzuki et al., 2023). In order to make (14) also satisfy Assumption 5-7, assumptions on $l$, $h$ and $r$ proposed in Nitanda et al. (2022) suffice. We only need to additionally specify the initialization without incurring stronger assumptions on $l$, $h$ and $r$. For instance, if we choose $r(x)=\|x\|^{2}/2$; there exists $\mathscr{L}>0$ such that the activation function satisfies $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (also proposed in Suzuki et al. (2023)); and the convex loss function $l$ is quadratic or satisfies $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (also proposed in Nitanda et al. (2022)), $F$ will meet Assumption 5 with $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp(-\frac{8\mathscr{L}}{d-2})$. Finally, if in addition we assume $l$ is $\sqrt{\mathscr{L}}$-Lipschitz and choose $\mu_{0}=\mathcal{N}(0,I_{2d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$, Assumptions 6 and 7 will be satisfied. We defer other examples and the related verification to Appendix B.
@@ 4.2 Density estimation
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (15) satisfies Assumptions 1-4 by choosing smooth and light-tailed kernel $k$, such as Gaussian radial basis function (RBF) kernel defined as $k(z,z^{\prime})\triangleq\exp(-\|z-z^{\prime}\|^{2}/2\sigma^{\prime 2})$ for $\sigma^{\prime}>0$. With Gaussian RBF kernel $k$ ($\sigma^{\prime}=\sigma$), we demonstrate that objective (15) also satisfies our Assumptions 5-7 if we choose $\lambda^{\prime}\lesssim 2\pi$, $r(x)=\|x\|^{2}/2$ and $\mu_{0}=\mathcal{N}(0,I_{2d})$. We defer the verification of our assumptions in the setting above to Appendix B.
@@ 5 Numerical experiments
$F$ satisfy Assumptions 1-7 with the choice of $l,\,h,\,r$, and thus we apply UNLA for minimizing the objective above. Note that the number of neurons in the first hidden layer is equivalent to the number of particles in NULA, and we choose $N\in\{256,512,1024,2048\}$.
@@ 6 Discussion
This paper (1) improves the convergence guarantees in Chen et al. (2023) with a refined Lyapunov analysis (Theorems 1 and 2); (2) discretizes the MULD and NULD with a scheme which results in smaller bias than the EM scheme; and (3) presents a novel discretization analysis of MULD and NULD. Our theoretical results suggest the mixing time of MULA and NULA might be superior to that of MLA and NLA, which is further validated by our experiments. We now note several directions for future potential developments. First, it is intriguing to see whether we can improve the number of particle guarantees for NULA to avoid the behavior of convergence instability in the low particle approximation regime. Second, we suspect convergence rates for MULD and NULD can be improved. Importantly, we obtain convergence rates for the MULA and NULA in TV distance, which do not share the same distance with the convergence rates of MULD, NULD, MLA and NLA. We hope to establish our results in the energy gap or KL divergence in the future. What’s more, our techniques on uniformly bounding the iterates of MULD and NULD combined with Assumptions 5-7 generates an additional $\mathscr{C}_{\textsf{LSI}}$ by Talagrand’s inequality, which is the reason why we do not improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ for MULA and NULA. We hope to explore whether it is possible to weaken those assumptions and refine the analysis of uniformly bounding the iterates to improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ in the mixing time and number of particles of MULA and NULA in our future work.
@@ Appendix A Helpful lemmas
Assume $F$ satisfies Assumptions 1-3. Then for every $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2d})$ we have
@@
Assume that $F$ satisfies Assumption 1 and there exists a measure $\mu_{*}\in\mathcal{P}(\mathbb{R}^{2d})$ that admits the proximal Gibbs distribution $\mu_{*}(x,v)\propto\exp\left(-\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2}\|v\|^{2}\right)$. Then for all $\mu^{N}\in\mathcal{P}(\mathbb{R}^{2dN})$, we have
@@
Let $(x_{t},v_{t})_{t\geq 0}$ and $(x_{t}^{i},v_{t}^{i})_{t\geq 0}$ respectively denote the iterates of the MULD and NULD. Assume that $h\lesssim\mathscr{L}^{-1/2}\wedge\gamma^{-1}$. Under Assumption 2 and Assumption 4, for $t\in[kh,(k+1)h]$, we have
@@
where the last inequality follows from Assumptions 2 and 4. Likewise for $V$:
@@
where the last inequality follows from Assumptions 2 and 4. Before applying matrix form of Grönwall’s inequality, let $c=c_{1}+c_{2}$ with $c_{2}=\left[\begin{array}[]{c}h\|v_{kh}\|\\ 0\end{array}\right]$,
@@
Let $(x_{t},v_{t})_{t\geq 0}$ denote the iterates of the MULD with $(x_{0},v_{0})\sim\mu_{0}=\mathcal{N}(0,I_{2d})$. Under Assumption 5 and Assumption 6, we have
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 3.222Assumption 3 states that the proximal Gibbs distribution satisfies the LSI. Note that $\mu_{*}$ also has the form of the proximal Gibbs distribution and thus satisfies LSI. The third inequality follows from Lemma 2. The fourth inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}(\mu_{t})<0$ along the MULD (Proof of Theorem 2.1 in Chen et al. (2023)) and the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}(\mu)$, we have $\mathcal{F}(\mu_{0})=F(\mu^{x}_{0})+\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)+\text{Ent}(\mu_{0})$. Since $(x_{0},v_{0})\sim\mathcal{N}(0,I_{2d})$, we have $\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)\lesssim d$ and
@@
By Assumption 6, we have $F(\mu_{0}^{x})\lesssim\mathscr{L}d$. By Assumption 5, we have $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@
Let ${(x^{i}_{t},v^{i}_{t})_{i=1}^{N}}$ denote the iterates of the NULD with $(x_{0}^{i},v_{0}^{i})\sim\mu_{0}^{i}=\mathcal{N}(0,I_{2d})$ for $i=1,...,N$ and $t\geq 0$. Under Assumption 5 and Assumption 7, we have
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 3. The third inequality follows from Lemma 4. The fourth inequality follows from Lemma 3 and the last inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{N}(\mu^{N}_{t})<0$ along the NULD (Proof of Theorem 2.2 in Chen et al. (2023)) and $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}^{N}(\mu^{N})$, we have $\mathcal{F}^{N}(\mu_{0}^{N})=\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})$. Similar to the proof of Lemma 7, since $(\textbf{x},\textbf{v})\sim\mathcal{N}^{\otimes N}(0,I_{2d})$, we have $\int\frac{1}{2}\|\textbf{v}\|^{2}\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim Nd$ and $|\text{Ent}(\mu_{0}^{N})|\lesssim Nd$. By Assumption 7 and Assumption 5, we also have $\int NF(\mu_{\textbf{x}})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim N\mathscr{L}d$ and $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@ B.1 Verification of Assumption 5
  • We will prove that Assumption 5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (such activation functions include tanh and sigmoid) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (such loss functions include logistic loss, Huber loss and log-cosh loss) or $l$ is quadratic. Since $r(x)=\|x\|^{2}/2$, we obtain
@@
  • We will prove that Assumption 5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$ (such activation functions include ReLU, GeLU, Softplus, SiLU) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$. Under these conditions, we have $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|\leq\mathscr{L}(1+\|x\|)$. Then, based on (24), we obtain
@@
We now prove that objective (15) satisfies Assumption 5 with Gaussian RBF kernel and $r(x)=\|x\|^{2}/2$. We choose $\sigma^{\prime}$ in Gaussian RBF kernel $k$ to be $\sigma$ for brevity. With the specific choice of $r$, we reformulate (15) as
@@ B.2 Verification of Assumption 6
Thus $F(\mu_{0})=\hat{\mathcal{M}}(\mu_{0})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}\lesssim\mathscr{L}+d,$ which satisfies Assumption 6.
B.3 Verification of Assumption 7
@@
Now we verify Assumption 7 for the example of density estimation. We consider the N-particle approximation of the objective (29) with $r(x)=\|x\|^{2}/2$ and initialization $\mu^{N}_{0}=\mathcal{N}^{\otimes N}(0,I_{d})$.
@@
which satisfies Assumption 7.
@@ Appendix C Continuous-time results
In this section, we give the explicit rate of Theorem 2.1 and Theorem 2.2 proposed by Chen et al. (2023) with a specific choice of parameters and then provide the detailed proof of Theorem 1 and Theorem 2 by reparameterizing $\gamma$.
C.1 Proof of Theorem 1
@@
which completes the proof of Theorem 1. Eq. 33 exhibits a faster rate than the rate of Eq. 32.
C.2 Proof of Theorem 2
@@
$\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*})\leq\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*})\leq(\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}t\right)+\frac{60\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}^{2}}$ (38)
which completes the proof of Theorem 2. The convergence rate exhibited in Eq. 38 is faster and incurs a smaller bias than the rate exhibited in Eq. 37.

### 01 Jan 2024 v2

@@ 4.1 Training mean-field neural networks
The objectives (14) satisfy Assumptions 2.1-2.4 for specific common choices $l$, $h$ and $r$ described in several works  (Nitanda et al., 2022; Chen et al., 2022, 2023; Suzuki et al., 2023). In order to make (14) also satisfy Assumptions 2.5-2.7, assumptions on $l$, $h$ and $r$ proposed in Nitanda et al. (2022) suffice. We only need to additionally specify the initialization without incurring stronger assumptions on $l$, $h$ and $r$. For instance, if we choose $r(x)=\|x\|^{2}/2$; there exists $\mathscr{L}>0$ such that the activation function satisfies $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (also proposed in Suzuki et al. (2023)); and the convex loss function $l$ is quadratic or satisfies $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (also proposed in Nitanda et al. (2022)), $F$ will meet Assumption 2.5 with $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp(-\frac{8\mathscr{L}}{d-2})$. Finally, if in addition we assume $l$ is $\sqrt{\mathscr{L}}$-Lipschitz and choose $\mu_{0}=\mathcal{N}(0,I_{2d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$, Assumptions 2.6 and 2.7 will be satisfied. We defer other examples and the related verification to Appendix B.
@@ 4.2 Density estimation
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (15) satisfies Assumptions 2.1-2.4 by choosing smooth and light-tailed kernel $k$, such as Gaussian radial basis function (RBF) kernel defined as $k(z,z^{\prime})\triangleq\exp(-\|z-z^{\prime}\|^{2}/2\sigma^{\prime 2})$ for $\sigma^{\prime}>0$. With Gaussian RBF kernel $k$ ($\sigma^{\prime}=\sigma$), we demonstrate that objective (15) also satisfies our Assumptions 2.5-2.7 if we choose $\lambda^{\prime}\lesssim 2\pi$, $r(x)=\|x\|^{2}/2$ and $\mu_{0}=\mathcal{N}(0,I_{2d})$. We defer the verification of our assumptions in the setting above to Appendix B.
@@ 5 Numerical experiments
$F$ satisfy Assumptions 2.1-2.7 with the choice of $l,\,h,\,r$, and thus we apply Algorithm 1 for minimizing the objective above. Note that the number of neurons in the first hidden layer is equivalent to the number of particles in NULA, and we choose $N\in\{256,512,1024,2048\}$.
@@ 6 Discussion
This paper (1) improves the convergence guarantees in Chen et al. (2023) with a refined Lyapunov analysis (Theorems 3.1 and 3.2); (2) discretizes the MULD and NULD with a scheme which results in smaller bias than the EM scheme; and (3) presents a novel discretization analysis of MULD and NULD. Our theoretical results suggest the mixing time of MULA and NULA might be superior to that of MLA and NLA, which is further validated by our experiments. We now note several directions for future potential developments. First, it is intriguing to see whether we can improve the number of particle guarantees for NULA to avoid the behavior of convergence instability in the low particle approximation regime. Second, we suspect convergence rates for MULD and NULD can be improved. Importantly, we obtain convergence rates for the MULA and NULA in TV distance, which do not share the same distance with the convergence rates of MULD, NULD, MLA and NLA. We hope to establish our results in the energy gap or KL divergence in the future. What’s more, our techniques on uniformly bounding the iterates of MULD and NULD combined with Assumptions 2.5-2.7 generates an additional $\mathscr{C}_{\textsf{LSI}}$ by Talagrand’s inequality, which is the reason why we do not improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ for MULA and NULA. We hope to explore whether it is possible to weaken those assumptions and refine the analysis of uniformly bounding the iterates to improve the dependence of $\mathscr{C}_{\textsf{LSI}}$ in the mixing time and number of particles of MULA and NULA.
@@ Appendix A Helpful lemmas
Assume $F$ satisfies Assumptions 2.1-2.3. Then for every $\mu\in\mathcal{P}_{2}(\mathbb{R}^{2d})$ we have
@@
Assume that $F$ satisfies Assumption 2.1 and there exists a measure $\mu_{*}\in\mathcal{P}(\mathbb{R}^{2d})$ that admits the proximal Gibbs distribution $\mu_{*}(x,v)\propto\exp\left(-\frac{\delta F}{\delta\mu}(\mu^{x}_{*},x)-\frac{1}{2}\|v\|^{2}\right)$. Then for all $\mu^{N}\in\mathcal{P}(\mathbb{R}^{2dN})$, we have
@@
Let $(x_{t},v_{t})_{t\geq 0}$ and $(x_{t}^{i},v_{t}^{i})_{t\geq 0}$ respectively denote the iterates of the MULD and NULD. Assume that $h\lesssim\mathscr{L}^{-1/2}\wedge\gamma^{-1}$. Under Assumption 2.2 and Assumption 2.4, for $t\in[kh,(k+1)h]$, we have
@@
where the last inequality follows from Assumptions 2.2 and 2.4. Likewise for $V$:
@@
where the last inequality follows from Assumptions 2.2 and 2.4. Before applying matrix form of Grönwall’s inequality, let $c=c_{1}+c_{2}$ with $c_{2}=\left[\begin{array}[]{c}h\|v_{kh}\|\\ 0\end{array}\right]$,
@@
Let $(x_{t},v_{t})_{t\geq 0}$ denote the iterates of the MULD with $(x_{0},v_{0})\sim\mu_{0}=\mathcal{N}(0,I_{2d})$. Under Assumption 2.5 and Assumption 2.6, we have
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3.222Assumption 2.3 states that the proximal Gibbs distribution satisfies the LSI. Note that $\mu_{*}$ also has the form of the proximal Gibbs distribution and thus satisfies LSI. The third inequality follows from Lemma 2. The fourth inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}(\mu_{t})<0$ along the MULD (Proof of Theorem 2.1 in Chen et al. (2023)) and the last inequality follows from the assumption that $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}(\mu)$, we have $\mathcal{F}(\mu_{0})=F(\mu^{x}_{0})+\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)+\text{Ent}(\mu_{0})$. Since $(x_{0},v_{0})\sim\mathcal{N}(0,I_{2d})$, we have $\int\frac{1}{2}\|v\|^{2}\mu_{0}(\mathrm{d}x\mathrm{d}v)\lesssim d$ and
@@
By Assumption 2.6, we have $F(\mu_{0}^{x})\lesssim\mathscr{L}d$. By Assumption 2.5, we have $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@
Let ${(x^{i}_{t},v^{i}_{t})_{i=1}^{N}}$ denote the iterates of the NULD with $(x_{0}^{i},v_{0}^{i})\sim\mu_{0}^{i}=\mathcal{N}(0,I_{2d})$ for $i=1,...,N$ and $t\geq 0$. Under Assumption 2.5 and Assumption 2.7, we have
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3. The third inequality follows from Lemma 4. The fourth inequality follows from Lemma 3 and the last inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{N}(\mu^{N}_{t})<0$ along the NULD (Proof of Theorem 2.2 in Chen et al. (2023)) and $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}^{N}(\mu^{N})$, we have $\mathcal{F}^{N}(\mu_{0}^{N})=\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})$. Similar to the proof of Lemma 7, since $(\textbf{x},\textbf{v})\sim\mathcal{N}^{\otimes N}(0,I_{2d})$, we have $\int\frac{1}{2}\|\textbf{v}\|^{2}\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim Nd$ and $|\text{Ent}(\mu_{0}^{N})|\lesssim Nd$. By Assumption 2.7 and Assumption 2.5, we also have $\int NF(\mu_{\textbf{x}})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim N\mathscr{L}d$ and $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@ Appendix B Verification of assumptions
  • We will prove that Assumption 2.5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}$ (such activation functions include tanh and sigmoid) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$ (such loss functions include logistic loss, Huber loss and log-cosh loss) or $l$ is quadratic. Since $r(x)=\|x\|^{2}/2$, we obtain
@@
  • We will prove that Assumption 2.5 holds if $r(x)=\|x\|^{2}/2$, $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$ (such activation functions include ReLU, GeLU, Softplus, SiLU) and $|\partial_{1}l|\leq\sqrt{\mathscr{L}}$. Under these conditions, we have $|\partial_{1}l(h(\mu;a_{i}),b_{i})h(x;a_{i})|\leq\mathscr{L}(1+\|x\|)$. Then, based on (24), we obtain
@@
We now prove that objective (15) satisfies Assumption 2.5 with Gaussian RBF kernel and $r(x)=\|x\|^{2}/2$. We choose $\sigma^{\prime}$ in Gaussian RBF kernel $k$ to be $\sigma$ for brevity. With the specific choice of $r$, we reformulate (15) as
@@
Thus $F(\mu_{0})=\hat{\mathcal{M}}(\mu_{0})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}\lesssim\mathscr{L}+d,$ which satisfies Assumption 2.6.
B.3 Verification of Assumption 2.7
@@
Now we verify Assumption 2.7 for the example of density estimation. We consider the N-particle approximation of the objective (29) with $r(x)=\|x\|^{2}/2$ and initialization $\mu^{N}_{0}=\mathcal{N}^{\otimes N}(0,I_{d})$.
@@
which satisfies Assumption 2.7.
@@ Appendix C Continuous-time results
In this section, we give the explicit rate of Theorem 2.1 and Theorem 2.2 proposed by Chen et al. (2023) with a specific choice of parameters and then provide the detailed proof of Theorem 3.1 and Theorem 3.2 by reparameterizing $\gamma$.
C.1 Proof of Theorem 3.1
@@
which completes the proof of Theorem 3.1. Eq. 33 exhibits a faster rate than the rate of Eq. 32.
C.2 Proof of Theorem 3.2
@@
$\mathcal{F}^{N}(\mu_{t}^{N})-N\mathcal{F}(\mu_{*})\leq\mathcal{E}^{N}(\mu_{t}^{N})-N\mathcal{E}(\mu_{*})\leq\mathcal{E}_{0}^{N}\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{6\sqrt{\mathscr{L}}}t\right)+\frac{60\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+\frac{36\mathscr{L}^{2}d}{\mathscr{C}_{\textsf{LSI}}^{2}}$ (38)
where $\mathcal{E}_{0}^{N}\triangleq\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}).$ This completes the proof of Theorem 3.2. The convergence rate exhibited in Eq. 38 is faster and incurs a smaller bias than the rate exhibited in Eq. 37.